the question is... write the equation of the line(in point-slope form) (-3,4)(5,0) and i completly forgot how to get the answer..i know how to graph it, but i dont remeber the formula...
Answer by tutorcecilia(2144)

(-3,4)(5,0)
m=(y2-y1)/(x2-x1) [slope format]
.
m=(0-4)/(5+3)=-4/8=-1/2 [find the slope]
.
y1-y2=m(x1-x2) [point-slope format]
y-0=-1/2(x-5) [plug-in the values for the point-slope format]